# Multi-target Range, Doppler and Angle estimation in MIMO-FMCW Radar with Limited Measurements

### v1

@@ IV-B Non-uniform recovery
We now examine the isotropy condition (21) for matrix $\mathbf{D}$ and non-uniform guarantees. Recall that any row of $\mathbf{D}$ is represented as $\mathbf{D}_{N_{T}N_{R}(i_{1}-1)+i_{2},:}=\mathbf{B}_{i_{1},:}\otimes\mathbf{C}_{i_{2},:}$ where $1\leq i_{1}\leq P$, $1\leq i_{2}\leq N_{T}N_{R}$. Now, any row of matrix $\mathbf{C}$ is $[\exp{(j\frac{\pi A}{\lambda}\sin{\phi_{1}}(\alpha_{n}+\beta_{m}))},\ldots,\exp{(j\frac{\pi A}{\lambda}\sin{\phi_{G_{\theta}}}(\alpha_{n}+\beta_{m}))}]$for some $1\leq n\leq N_{T}$ and $1\leq m\leq N_{R}$. Hence, rows of $\mathbf{C}$ and, consequently, $\mathbf{D}$ are generally not independent. Non-uniform recovery for such cases has been addressed in [53]. Here, we first provide conditions to satisfy the isotropy property in Theorem 3. The non-uniform recovery guarantees then follow in Theorem 4.
@@
Consider the random sparse chirps and SLA setup of Theorem 1. Also, consider a given $K$-sparse signal $\mathbf{z}$ measured by (19) with $\|\mathbf{w}\|_{2}\leq\sigma$. We assume that the distributions $\mathcal{P}_{p}$, $\mathcal{P}_{\alpha}$ and $\mathcal{P}_{\beta}$ and uniform grids $\rho_{1:G_{D}}$ and $\phi_{1:G_{\theta}}$ satisfy (32). Consider some $10<\epsilon<$ 1 such that
@@
Using the isotropy property, the non-uniform recovery of $K$-targets can be guaranteed by generalizing [53, Theorem 2.1] for the vectorized model (19). Particularly, we consider the exact $K$-sparse signal $\mathbf{z}$ and substitute the following: (a) number of rows of measurement matrix by $PN_{T}N_{R}$, (b) sparsity by $K$, and (c) number of columns of measurement matrix by $G_{D}G_{\theta}$. ∎
@@ IV-C Practical example
In Proposition 5, if $\phi_{1:G_{\theta}}$ covers the entire $[-\pi/2,\pi/2]$ interval, then $G_{\theta}=(A/\lambda)+1$ where $A/\lambda$ is the normalized array aperture. However, in practice, the radar focuses on a specific angular sector in each CPI. Similarly, denote $v_{max}=\lambda/4T_{c}$ as the maximum target velocity that can be detected by the given radar system. If $\rho_{1:G_{D}}$ covers $[-v_{max},v_{max}]$ interval, then $G_{D}=P_{max}+1$. Interestingly, the velocity resolution is then $\lambda/(2P_{max}T_{c})$, same as the traditional radar transmitting $P_{max}$ chirps. However, the proposed framework achieves this resolution by transmitting only $P<P_{max}$ chirps.
@@ V Simulation results
We compare the detection and estimation performance of our proposed methods with the classical-DFT and MUSIC techniques. Section V-A considers various range estimation methods while Section V-B focuses on the overall performance and time-complexity of various methods. The simulation parameters are listed in Table I. For SLA, $\alpha_{n},\beta_{m}\sim\mathcal{U}[-1/2,1/2]$ for all $1\leq n\leq N_{T}$ and $1\leq m\leq N_{R}$. For ULA, we consider the MIMO array setup of [12] wherein two transmitters are placed on either side of the array with $\lambda$ spacing, and the receivers are placed between them with $0.5\lambda$ and $0.25\lambda$ spacings with one another and the transmitters, respectively. This arrangement results in $20$ unique virtual array channels. Similarly, for the sparsely transmitted chirps, $\zeta_{1\leq p\leq P}$ are chosen uniformly and independently from $\{0,1,\ldots,P_{max}-1\}$ without replacement, while the classical FMCW setup transmits all $P_{max}$ chirps. The target gains are generated as $a_{k}=\exp{(j\psi_{k})}$ with $\psi_{k}\sim\mathcal{U}[0,2\pi)$ (without any path loss) while the noise $w_{n,m,p}[t]\sim\mathcal{CN}(\mathbf{0},\sigma^{2}I)$, i.i.d. across all chirps and virtual array channels. Note that the SNR is $-10\log_{10}(\sigma^{2})$[31].
TABLE I: Simulation Parameters
@@
Number of antenna elements in ULA $N_{T}=4$, $N_{R}=8$
@@
V-A DFT-focusing and Range-OMP
Consider three close range targets with ranges $R_{1}=48.8$m, $R_{2}=50.0$m and $R_{3}=51.2$m for the full measurements radar setup. For the parameters given in Table I, the range resolution for the DFT processing is $0.6$m. In range-OMP, we choose a uniform grid $\omega_{1\leq g\leq G_{R}}$ covering ranges $1.2$m to $120$m with resolution $0.12$m. Both the binary and coherent integration process all $20\times 32$ measurements $y_{n,m,p}$, but binary integration applies a $1/3$-rd majority rule to filter valid target ranges. Contrarily, range-OMP considers only $y_{1,1,1}$ channel.
Fig. 2 shows the normalized range channels for DFT (with coherent and binary integration) and range-OMP. Fig. 2a & b indicate that only two targets are detected in coherent processing and Channel-A of binary processing. However, in the case of binary integration, the missed target is detected at Channel-B in Fig. 2c. On the contrary, the range-OMP in Fig. 2d recovers all three targets at a higher resolution using only $y_{1,1,1}$. Hence, range-OMP provides better estimates at lower complexity than DFT. Note that binary integration enhances performance only at low SNRs and exhibits the same performance as the coherent one at higher SNRs. Similarly, range-OMP followed by binary integration can be used to detect ranges from all $\{y_{n,m,p}\}$ measurements, resulting in enhanced detection at low SNRs but with increased computational costs. Throughout our experiments, we consider the range-OMP for $y_{1,1,1}$ only.
@@ V-B Range, velocity and AOA estimation
We consider $K=5$ targets with uniformly drawn ranges $R_{k}\in[20\;m,120\;m]$, velocities $\nu_{k}\in[-78\;m/s,78\;m/s]$ and AOAs $\theta_{k}\in[-20^{\circ},20^{\circ}]$. We compare our proposed CS-based methods for SLA and sparse chirps with the conventional full-measurement radar systems in terms of detection probabilities, estimation errors, ROC, and run-time complexity. Table-II describes various methods considered for comparison. For classical-DFT, the range, velocity, and angular resolutions are $0.6$m, $4.88\;m/s$, and $7^{\circ}$, respectively, for the parameters given in Table I. In range-OMP, a grid of resolution $0.12$m is chosen as in Section V-V-A. Similarly, the velocity grid $\rho_{1\leq g\leq G_{D}}$ is constructed using $200$ uniform points in $[-78\;m/s,78\;m/s]$ with resolution $0.78\;m/s$. The angular grid $\phi_{1\leq g\leq G_{\theta}}$ is a uniform grid (in $\sin\theta$ domain) of $50$ points in $[-30^{\circ},30^{\circ}]$ resulting in an approximate angular resolution of $2^{\circ}$. The CS-based AOA estimation method proposed in [31] (for pulsed-wave radars) assumes a known sparsity level and uses prior knowledge of the actual number of targets $K$ in the CS algorithms. On the contrary, in 1D-OMP (19) and 2D-OMP (18) methods, we consider a sparsity level of $K_{max}=20$ to terminate the algorithms, while BP and LASSO consider $100$ iterations. A threshold detection is then used to obtain the target velocities and AOAs from the recovered signal.
TABLE II: Various parameter estimation methods
Method Radar Description
Classical-DFT ULA with DFT in range, Doppler and angular domains
$P_{max}$ chirps
MUSIC ULA with DFT in range and 2D-MUSIC with
@@
$P_{max}$ chirps spatial smoothing[12] in Doppler-angle domain
@@
OMP+Binary SLA with DFT focusing with binary integration for range
sparse $P$ chirps and 1D-OMP with (19) for velocity & AOA
OMP+Range-OMP SLA with Range-OMP and
sparse $P$ chirps 1D-OMP with (19) for velocity & AOA
2D-OMP SLA with DFT-with-binary for range and
sparse $P$ chirps 2D-OMP with (18) for velocity & AOA
BP SLA with DFT-with-binary for range and
sparse $P$ chirps 1D-BP with (19) for velocity & AOA
LASSO SLA with DFT-with-binary for range and
sparse $P$ chirps 1D-LASSO with (19) for velocity & AOA
@@
Since different methods have different range, Doppler and angular resolutions, we investigate the target detection and parameter estimation performances separately. To this end, we define a detected target as ‘hit’ if the estimated range, velocity, and AOA from a given method are within the corresponding resolution from the true target parameters. Otherwise, it is a false alarm. We compare the hit rates of various methods to demonstrate target detection performance while different thresholds are set to maintain a constant false-alarm rate at all SNRs. For parameter estimation, we compute root-mean-squared estimation error (RMSE) for the targets classified as hits. Unless mentioned otherwise, the rates and RMSEs are averaged over $300$ independent runs.
@@
1) Comparison with classical-DFT & MUSIC: Fig. 4 compares the false-alarm and hit rates of various methods over $-25$dB to $40$dB SNRs. The corresponding range, velocity, and angle RMSEs are shown in Fig. 3. Fig. 4 demonstrates that our proposed ‘OMP+Binary’ method achieves similar false-alarm and hit rates as the classical-DFT and MUSIC for all SNRs, but with only half the number of antenna elements and less than half transmitted chirps. In particular, OMP’s sparse radar setup considers $6$ elements ($2$ transmitters and $4$ receivers), while the classical-DFT and MUSIC consider a ULA of $12$ elements ($4$ transmitters and $8$ receivers). Similarly, the conventional setup transmits $32$ chirps in a CPI, while in the sparse scenario, only $10$ are transmitted randomly. Range-OMP provides similar performance only at high SNRs, but at lower computational costs than DFT-focusing as indicated in Table III.
In Fig. 3a, range-OMP, owing to its fine-resolution grid, exhibits significantly lower range estimation error than the DFT approach employed in other methods. Similarly, in Fig. 3b-c, the fine-resolution velocity and angle grids result in lower RMSEs than classical-DFT, but comparable to MUSIC. However, MUSIC, being a subspace-based approach, provides higher resolution with full measurements and higher computational costs.
2) Variation with number of targets: Fig. 5 shows the hit rate for different number of targets $K$ at $30$dB and $-20$dB SNRs, for OMP+Binary, classical-DFT and MUSIC methods. The false-alarm rate is kept constant at $0.05$. At high SNR, our proposed OMP method surpasses MUSIC with a hit rate greater than $0.95$, like classical DFT. As $K$ increases, MUSIC’s hit rate decreases because of the fixed smoothing parameters and higher correlation among targets. On the other hand, OMP and classical-DFT maintain a near-constant hit rate. At very low SNR ($-20$dB), both OMP and MUSIC’s hit rate degrades rapidly with $K$, unlike classical-DFT. While at $30$dB SNR, our proposed method surpasses MUSIC for all $K$, MUSIC exceeds at $-20$dB. OMP’s recovery depends heavily on the measurement matrices and hence, the velocity and angular grids. A coarser grid with small correlation among dictionary elements can enhance the hit rate at low SNRs, trading off resolution.
@@
Figure 6: ROC curves for OMP+Binary, classical-DFT, and MUSIC methods at $30$dB and $-20$dB SNRs.
3) ROCs: Fig. 6 shows the ROCs for OMP+Binary, classical-DFT and MUSIC methods at $30$dB and $-20$dB SNRs. All methods exhibit near-ideal ROCs at both low and high SNRs. Unlike classical-DFT, the detection performance of MUSIC and OMP degrades slightly at $-20$dB SNR, with OMP’s ROC being below MUSIC. However, as mentioned earlier, MUSIC processes measurements from a $4\times 8$ ULA and all $32$ transmitted chirps, while our method considers only a $2\times 4$ SLA and $10$ chirps in a CPI. Furthermore, from Table III, MUSIC has approximately $2.5$ times longer run-time than OMP.
4) Comparison of different CS methods: Fig. 7 illustrates the false-alarm and hit rates of 1-D OMP, BP, and LASSO as well as 2D-OMP methods. Because of the same velocity and angular grids, the RMSEs of these methods are the same and, hence, omitted here. All algorithms exhibit similar false-alarm and hit rates for high SNRs ($\geq 0$dB). On the other hand, at low SNRs ($-20$dB to 00dB), LASSO achieves the same hit rate as other methods at the lowest false-alarm rate and hence, is more robust to noise. However, these methods differ in run-time complexity as elaborated in the following.
@@
TABLE III: Run-time for different methods (at $15$dB SNR)
@@
5) Time complexity: Table III provides the run-time (for one run) for various methods listed in Table II. As expected, MUSIC exhibits the longest run-time even though it provides high resolutions. The classical-DFT and OMP+Binary methods demonstrate similar time complexity, but OMP is more accurate, as indicated in Fig. 3. Range-OMP further reduces the run-time of our method because it considers only measurement $y_{1,1,1}$ for range detection, see Section V-A. Among different CS methods, 2D-OMP significantly outperforms the 1D OMP, BP, and LASSO algorithms. Overall, our CS-based methods, with 2D-OMP being the most efficient, achieve similar hit rates but higher resolutions than classical methods, even with limited measurements. Our range-OMP further outperforms DFT-based range processing in terms of both resolution and computational costs.

### v2

@@ IV-B Non-uniform recovery
We now examine the isotropy condition (22) for matrix $\mathbf{D}$ and non-uniform guarantees. Recall that any row of $\mathbf{D}$ is represented as $\mathbf{D}_{N_{T}N_{R}(i_{1}-1)+i_{2},:}=\mathbf{B}_{i_{1},:}\otimes\mathbf{C}_{i_{2},:}$ where $1\leq i_{1}\leq P$, $1\leq i_{2}\leq N_{T}N_{R}$. Now, any row of matrix $\mathbf{C}$ is $\exp{(j\frac{\pi A}{\lambda}\sin{\phi_{1}}(\alpha_{n}+\beta_{m}))},\ldots,\exp{(j\frac{\pi A}{\lambda}\sin{\phi_{G_{\theta}}}(\alpha_{n}+\beta_{m}))}]$for some $1\leq n\leq N_{T}$ and $1\leq m\leq N_{R}$. Hence, rows of $\mathbf{C}$ and, consequently, $\mathbf{D}$ are generally not independent. Non-uniform recovery for such cases has been addressed in [54]. Here, we first provide conditions to satisfy the isotropy property in Theorem 3. The non-uniform recovery guarantees then follow in Theorem 4.
@@
Consider the random sparse chirps and SLA setup of Theorem 1. Also, consider a given $K$-sparse signal $\mathbf{z}$ measured by (20) with $\|\mathbf{w}\|_{2}\leq\sigma$. We assume that the distributions $\mathcal{P}_{p}$, $\mathcal{P}_{\alpha}$ and $\mathcal{P}_{\beta}$ and uniform grids $\rho_{1:G_{D}}$ and $\phi_{1:G_{\theta}}$ satisfy (33). Consider some $0<\epsilon<1$ such that
@@
Using the isotropy property, the non-uniform recovery of $K$-targets can be guaranteed by generalizing [54, Theorem 2.1] for the vectorized model (20). Particularly, we consider the exact $K$-sparse signal $\mathbf{z}$ and substitute the following: (a) number of rows of measurement matrix by $PN_{T}N_{R}$, (b) sparsity by $K$, and (c) number of columns of measurement matrix by $G_{D}G_{\theta}$. ∎
@@
Intuitively, bounds (30)-(32) from Theorem 2 and bound (34) from Theorem 4 provide the sufficient number of measurements necessary to control the off-diagonal elements of the measurement matrix $\mathbf{D}$, thereby ensuring high-probability recovery. In (28)-(29), the coherence $\mu$, also viewed as the peak sidelobe of the array pattern[32], provides a bound on the likelihood of this peak exceeding the threshold $\vartheta$. Meanwhile, the isotropy condition $\mathbb{E}[\mathbf{D}_{l,:}^{H}\mathbf{D}_{l,:}]=\mathbf{I}$ resembles the aperture condition in [54] and serves to regulate the variance of the non-diagonal elements of the matrix $\mathbf{Q}_{D}$. Finally, conditions (26)-(27) and (33) imply that the grid points should be located at the zeros of the corresponding characteristic functions. We also note that Theorems 1-4 provide sufficient—though not necessary—conditions for ensuring high-probability recovery. In practice, the target’s velocity and AOA in (20) are estimated from the support of the sparse vector $\mathbf{z}$, rather than from the exact values of its non-zero entries. Consequently, in our experiments in Section V, the proposed methods can match the detection performance of classical-DFT and MUSIC techniques, even when using only half the number of chirps and antenna elements.
@@ IV-C Practical example
In Proposition 5, if $\phi_{1:G_{\theta}}$ covers the entire $\pi/2,\pi/2]$ interval, then $G_{\theta}=(A/\lambda)+1$ where $A/\lambda$ is the normalized array aperture. However, in practice, the radar focuses on a specific angular sector in each CPI. Similarly, denote $v_{max}=\lambda/4T_{c}$ as the maximum target velocity that can be detected by the given radar system. If $\rho_{1:G_{D}}$ covers $v_{max},v_{max}]$ interval, then $G_{D}=P_{max}+1$. Interestingly, the velocity resolution is then $\lambda/(2P_{max}T_{c})$, same as the traditional radar transmitting $P_{max}$ chirps. However, the proposed framework achieves this resolution by transmitting only $P<P_{max}$ chirps.
@@ V Simulation results
We compare the detection and estimation performance of our proposed methods with the classical-DFT and MUSIC techniques. Section V-A considers various range estimation methods while Section V-B focuses on the overall performance and time-complexity of various methods. The simulation parameters are listed in Table III. For SLA, $\alpha_{n},\beta_{m}\sim\mathcal{U}[-1/2,1/2]$ for all $1\leq n\leq N_{T}$ and $1\leq m\leq N_{R}$. For ULA, we consider the MIMO array setup of [12] wherein two transmitters are placed on either side of the array with $\lambda$ spacing, and the receivers are placed between them with $0.5\lambda$ and $0.25\lambda$ spacings with one another and the transmitters, respectively. This arrangement results in $20$ unique virtual array channels. Similarly, for the sparsely transmitted chirps, $\zeta_{1\leq p\leq P}$ are chosen uniformly and independently from $\{0,1,\ldots,P_{max}-1\}$ without replacement, while the classical FMCW setup transmits all $P_{max}$ chirps. The target gains are generated as $a_{k}=\exp{(j\psi_{k})}$ with $\psi_{k}\sim\mathcal{U}[0,2\pi)$ (without any path loss) while the noise $w_{n,m,p}[t]\sim\mathcal{CN}(\mathbf{0},\sigma^{2}I)$, i.i.d. across all chirps and virtual array channels. Note that the SNR is $10\log_{10}(\sigma^{2})$[32]. In all experiments, both the random SLA and ULA are assumed to be perfectly calibrated. In Appendix F, we demonstrate that the proposed methods with random SLAs exhibit similar sensitivity to phase and delay calibration errors as the subspace-based MUSIC approach with ULA, which is higher than the classical-DFT.
V-A DFT-focusing and Range-OMP
Consider three close range targets with ranges $R_{1}=48.8$m, $R_{2}=50.0$m and $R_{3}=51.2$m for the full measurements radar setup. For the parameters given in Table III, the range resolution for the DFT processing is 0.60.6m. In Range-OMP, we choose a uniform grid $\omega_{1\leq g\leq G_{R}}$ covering ranges 1.21.2m to 120120m with resolution 0.120.12m. Both the binary and coherent integration process all $20\times 32$ measurements $y_{n,m,p}$, but binary integration applies a $1/3$-rd majority rule to filter valid target ranges. Contrarily, Range-OMP considers only $y_{1,1,1}$ channel.
Fig. 3 shows the normalized range channels for DFT (with coherent and binary integration) and Range-OMP at 20-20dB SNR. Fig. 3a & b indicate that only two targets are detected in coherent processing and Channel-A of binary processing. However, in the case of binary integration, the missed target is detected at Channel-B in Fig. 3c. On the contrary, the Range-OMP in Fig. 3d recovers all three targets at a higher resolution using only $y_{1,1,1}$. Hence, Range-OMP provides better estimates at lower complexity than DFT. Note that binary integration enhances performance only at low SNRs and exhibits the same performance as the coherent one at higher SNRs. Similarly, Range-OMP followed by binary integration can be used to detect ranges from all $\{y_{n,m,p}\}$ measurements, resulting in enhanced detection at low SNRs but with increased computational costs. Throughout our experiments, we consider the Range-OMP for $y_{1,1,1}$ only.
TABLE III: Simulation Parameters
@@ V-A DFT-focusing and Range-OMP
Number of antenna elements in ULA $N_{T,f}=4$, $N_{R,f}=8$
@@ V-B Range, velocity and AOA estimation
We consider $K=5$ targets with uniformly drawn ranges $R_{k}\in[20\;m,120\;m]$, velocities $\nu_{k}\in[-78\;m/s,78\;m/s]$ and AOAs $\theta_{k}\in[-20^{\circ},20^{\circ}]$. We compare our proposed CS-based methods for SLA and sparse chirps with the conventional full-measurement radar systems in terms of detection probabilities, estimation errors, ROC, and run-time complexity. Table-IV describes various methods considered for comparison. For classical-DFT, the range, velocity, and angular resolutions are 0.60.6m, 4.88m/s4.88\;m/s, and $7^{\circ}$, respectively, for the parameters given in Table III. In Range-OMP, a grid of resolution 0.120.12m is chosen as in Section V-A. Similarly, the velocity grid $\rho_{1\leq g\leq G_{D}}$ is constructed using $200$ uniform points in [78m/s,78m/s][-78\;m/s,78\;m/s] with resolution 0.78m/s0.78\;m/s. The angular grid $\phi_{1\leq g\leq G_{\theta}}$ is a uniform grid (in $\sin\theta$ domain) of $50$ points in $30^{\circ},30^{\circ}]$ resulting in an approximate angular resolution of $2^{\circ}$. The CS-based AOA estimation method proposed in [32] (for pulsed-wave radars) assumes a known sparsity level and uses prior knowledge of the actual number of targets $K$ in the CS algorithms. On the contrary, in Range-OMP (14), 1D-OMP (20), and 2D-OMP (19) methods, we considered a sparsity level of $K_{max}=20$ to terminate the algorithms, while BP and LASSO considered $100$ iterations and an optimization tolerance of $10^{-4}$. The regularization parameter for the LASSO method was set to $0.6$. A threshold detection is then used to obtain the target parameters from the recovered signal. Note that, consistent with standard CS algorithms, the computational complexity of these methods is also determined by the chosen sparsity level $K_{max}$ and the number of iterations, as detailed in Table II. Additionally, while these parameters influence detection performance, their impact on estimation error is minimal. The estimation accuracy is instead primarily determined by the resolution of the corresponding grids.
TABLE IV: Various parameter estimation methods
Method Radar setup Description
Classical-DFT ULA with $P_{max}$ chirps DFT in range, Doppler and angular domains
@@
MUSIC ULA with $P_{max}$ chirps DFT in range and 2D-MUSIC with spatial smoothing[12] in Doppler-angle domain
@@
OMP+Binary SLA with sparse $P$ chirps DFT focusing with binary integration for range and 1D-OMP with (20) for velocity & AOA
OMP+Range-OMP SLA with sparse $P$ chirps Range-OMP and 1D-OMP with (20) for velocity & AOA
2D-OMP SLA with sparse $P$ chirps DFT-with-binary for range and 2D-OMP with (19) for velocity & AOA
BP SLA with sparse $P$ chirps DFT-with-binary for range and 1D-BP with (20) for velocity & AOA
LASSO SLA with sparse $P$ chirps DFT-with-binary for range and 1D-LASSO with (20) for velocity & AOA
@@
Since different methods have different range, Doppler and angular resolutions, we investigate the target detection and parameter estimation performances separately. To this end, we define a detected target as ‘hit’ if the estimated range, velocity, and AOA from a given method are within the corresponding resolution from the true target parameters. Otherwise, it is a false alarm. We compare the hit rates of various methods to demonstrate target detection performance while different thresholds are set to maintain a near-constant false-alarm rate at all SNRs. For parameter estimation, we compute root-mean-squared estimation error (RMSE) for the targets classified as hits. Unless mentioned otherwise, the rates and RMSEs are averaged over $300$ independent runs.
@@
1) Comparison with classical-DFT & MUSIC: Fig. 5 compares the false-alarm and hit rates of various methods over 25-25dB to 4040dB SNRs. The corresponding range, velocity, and angle RMSEs are shown in Fig. 4. Fig. 5 demonstrates that our proposed ‘OMP+Binary’ method achieves similar false-alarm and hit rates as the classical-DFT and MUSIC for SNRs above 20-20 dB, but with only half the number of antenna elements and less than half transmitted chirps. Note that the performance degradation at very low SNRs, compared to uniform measurement systems, is a known limitation of radar frameworks relying on sparse signal processing [29, 26, 58]. The ULA-based classical-DFT and MUSIC achieve a hit rate of $0.9$ even at 25-25 dB, benefiting from the larger number of measurements. However, practical MIMO-FMCW radars usually operate in the 0200-20 dB SNR regime at the receiver’s input. In particular, OMP’s sparse radar setup considers $6$ elements ($2$ transmitters and $4$ receivers), while the classical-DFT and MUSIC consider a ULA of $12$ elements ($4$ transmitters and $8$ receivers). Similarly, the conventional setup transmits $32$ chirps in a CPI, while in the sparse scenario, only $10$ are transmitted randomly. Range-OMP provides similar performance only at high SNRs, but at lower computational costs than DFT-focusing as indicated in Table V.
In Fig. 4a, Range-OMP, owing to its fine-resolution grid, exhibits significantly lower range estimation error than the DFT approach employed in other methods. Similarly, in Fig. 4b-c, the fine-resolution velocity and angle grids result in lower RMSEs than classical-DFT, but comparable to MUSIC. However, MUSIC, being a subspace-based approach, provides higher resolution with full measurements and higher computational costs. Note that while SNR primarily influences the hit and false alarm rates of a method, the estimation error is largely determined by the resolution associated with the method’s parameter grids.
@@
Figure 7: ROC curves for OMP+Binary, classical-DFT, and MUSIC methods at 3030dB and 20-20dB SNRs.
2) Variation with number of targets: Fig. 6 shows the hit rate for different number of targets $K$ at 3030dB and 20-20dB SNRs, for OMP+Binary, classical-DFT and MUSIC methods. The false-alarm rate is kept constant at $0.05$. At high SNR, our proposed OMP method surpasses MUSIC with a hit rate greater than $0.95$, like classical DFT. As $K$ increases, MUSIC’s hit rate decreases because of the fixed smoothing parameters and higher correlation among targets. On the other hand, OMP and classical-DFT maintain a near-constant hit rate. At very low SNR (20-20dB), both OMP and MUSIC’s hit rate degrades rapidly with $K$, unlike classical-DFT. While at 3030dB SNR, our proposed method surpasses MUSIC for all $K$, MUSIC exceeds at 20-20dB. OMP’s recovery depends heavily on the measurement matrices and hence, the velocity and angular grids. A coarser grid with small correlation among dictionary elements can enhance the hit rate at low SNRs, trading off resolution.
@@
TABLE V: Run-time for different methods (at 1515dB SNR)
@@
3) ROCs: Fig. 7 shows the ROCs for OMP+Binary, classical-DFT and MUSIC methods at 3030dB and 20-20dB SNRs. All methods exhibit near-ideal ROCs at both low and high SNRs. Unlike classical-DFT, the detection performance of MUSIC and OMP degrades slightly at 20-20dB SNR, with OMP’s ROC being below MUSIC. However, as mentioned earlier, MUSIC processes measurements from a $4\times 8$ ULA and all $32$ transmitted chirps, while our method considers only a $2\times 4$ SLA and $10$ chirps in a CPI. Also, from Table V, MUSIC has approximately $2.5$ times longer run-time than OMP.
4) Comparison of different CS methods: Fig. 8 illustrates the false-alarm and hit rates of 1-D OMP, BP, and LASSO as well as 2D-OMP methods. Because of the same velocity and angular grids, the RMSEs of these methods are the same and, hence, omitted here. All algorithms exhibit similar false-alarm and hit rates for high SNRs ($\geq 0$dB). On the other hand, at low SNRs (20-20dB to 0dB), LASSO achieves the same hit rate as other methods at the lowest false-alarm rate and hence, is more robust to noise. However, these methods differ in run-time complexity as elaborated in the following.
5) Time complexity: Table V provides the run-time (for one run) [^2] We conducted our experiments on Dell Latitude 3410 with Intel Core i7 $10^{th}$ generation CPU @ 1.80 GHz clock speed and 8 GB RAM, using MATLAB (R2023a) on Windows 11. for various methods listed in Table IV. As expected, MUSIC exhibits the longest run-time even though it provides high resolutions. The classical-DFT and OMP+Binary methods demonstrate similar time complexity, but OMP is more accurate, as indicated in Fig. 4. Range-OMP further reduces the run-time of our method because it considers only measurement $y_{1,1,1}$ for range detection; see Section V-A. Among different CS methods, 2D-OMP significantly outperforms the 1D OMP, BP, and LASSO algorithms. Overall, our CS-based methods, with 2D-OMP being the most efficient, achieve similar hit rates but higher resolutions than classical methods, even with limited measurements. Owing to the reduced number of measurements, our OMP methods are also faster than classical-DFT. Our Range-OMP further outperforms DFT-based range processing in terms of both resolution and computational costs.